OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 -x - y = -7 Answer y - 7 y + 7 -y - 7 -y + 7

OpenStudy (rogue):

-x - y = -7 x + y = 7 x = 7 - y = -y + 7